# Robust Data Pruning: Uncovering and Overcoming Implicit Bias

### v1

@@ Abstract
In the era of exceptionally data-hungry models, careful selection of the training data is essential to mitigate the extensive costs of deep learning. Data pruning offers a solution by removing redundant or uninformative samples from the dataset, which yields faster convergence and improved neural scaling laws. However, little is known about its impact on classification bias of the trained models. We conduct the first systematic study of this effect and reveal that existing data pruning algorithms can produce highly biased classifiers. At the same time, we argue that random data pruning with appropriate class ratios has potential to improve the worst-class performance. We propose a “fairness-aware” approach to pruning and empirically demonstrate its performance on standard computer vision benchmarks. In sharp contrast to existing algorithms, our proposed method continues improving robustness at a tolerable drop of average performance as we prune more from the datasets. We present theoretical analysis of the classification risk in a mixture of Gaussians to further motivate our algorithm and support our findings.
@@ 1 Introduction
The ever-increasing state-of-the-art performance of deep learning models requires exponentially larger volumes of training data according to neural scaling laws (Hestness et al., 2017; Kaplan et al., 2020; Rosenfeld et al., 2020; Gordon et al., 2021). However, not all collected data is equally important for learning as it contains noisy, repetitive, or uninformative samples. A recent research thread on data pruning is concerned with removing those unnecessary data, resulting in improved convergence speed, neural scaling factors, and resource efficiency (Toneva et al., 2019; Paul et al., 2021; He et al., 2023). These methods design scoring mechanisms to assess the utility of each sample, often measured by its difficulty or uncertainty as approximated during a preliminary training round, that guides pruning. Sorscher et al. (2022) report that selecting high-quality data using these techniques can trace a Pareto optimal frontier, beating the notorious power scaling laws.
@@
The modus operandi of data pruning resembles algorithms that mitigate distributional bias—a well-established issue in AI systems concerning the performance disparity across protected groups of the population (e.g., race or gender) or classes (Dwork et al., 2012; Hardt et al., 2016). The early approaches in this domain improve the worst-class performance of imbalanced datasets by subsampling the majority classes (Barua et al., 2012; Cui et al., 2019; Tan et al., 2020). Data pruning extends this strategy by designing finer pruning criteria and working for balanced datasets, too. More recent developments apply importance weighting to emphasize under-represented or under-performing groups or classes during optimization (Sinha et al., 2022; Sagawa* et al., 2020; Liu et al., 2021; Idrissi et al., 2022; Wang et al., 2023; Lukasik et al., 2022; Chen et al., 2017). Data pruning adheres to this paradigm as well by effectively assigning zero weights to the removed training samples. These parallels indicate that data pruning has potential to reduce classification bias in deep learning models, all while offering greater resource and memory efficiency. In Section 2, we conduct the first systematic evaluation of various pruning algorithms with respect to distributional robustness and find that this potential remains largely unrealized. For example, we find that Dynamic Uncertainty by He et al. (2023) achieves superior average test performance on CIFAR-100 with VGG-19 but fails miserably in terms of worst-class accuracy. Thus, it is imperative to benchmark pruning algorithms using a more comprehensive suite of performance metrics that reflect classification bias, and to develop solutions that address fairness directly.
@@
In contrast to data pruning, many prior works on distributionally robust optimization compute difficulty-based importance scores at the level of groups or classes and not for individual training samples. For pruning, this would correspond to selecting appropriate class ratios but subsampling randomly within each class. To imitate this behavior, we propose to select class proportions based on the corresponding class-wise error rates computed on a hold-out validation set after a preliminary training on the full dataset, and call this procedure Metric-based Quotas (MetriQ). While these quotas already improve robustness of the existing pruning algorithms, they particularly shine when applied together with random pruning, substantially reducing the classification bias of models trained on the full dataset while offering an enhanced data efficiency as a bonus. Our pruning protocol even achieves higher robustness than one cost-sensitive learning approach that has access to the full dataset (Sinha et al., 2022), at no degradation of the average performance for low enough pruning ratios. A theoretical analysis of linear decision rules for a mixture of two univariate Gaussians presented in Section 4 illustrates why simple random subsampling according to MetriQ yields better worst-class performance compared to existing, finer pruning algorithms.
@@
  • In Section 2, using a standard computer vision testbed, we conduct the first comprehensive evaluation of existing data pruning algorithms through the lens of fairness;
  • In Section 3, we propose a random pruning procedure with error-based class ratios coined MetriQ, and verify its effectiveness in drastically reducing the classification bias;
  • In Section 4, we back our empirical work with theoretical analysis of a toy Gaussian mixture model to illustrate the fundamental principles that contribute to the success of MetriQ.
2 Data Pruning & Fairness
As deep learning models become more data-hungry, researchers and practitioners are focusing extensively on improving data efficiency. Indeed, there is much room for improvement: it is generally accepted that large-scale datasets have many redundant, noisy, or uninformative samples that may even have a negative effect on learning, if any (He et al., 2023; Chaudhuri et al., 2023). Thus, it is imperative to carefully curate a subset of all available data for learning purposes. The corresponding literature is exceptionally rich, with a few fruitful and relevant research threads. Dataset distillation replaces the original examples with synthetically generated samples that bear compressed, albeit not as much interpretable, training signal (Sucholutsky and Schonlau, 2021; Cazenavette et al., 2022; Such et al., 2020; Zhao and Bilen, 2023; Nguyen et al., 2021; Feng et al., 2024). Coreset methods select representative samples that jointly capture the entire data manifold (Sener and Savarese, 2018; Guo et al., 2022; Zheng et al., 2023; Agarwal et al., 2020; Mirzasoleiman et al., 2020; Welling, 2009); they yield weak generalization guarantees for non-convex problems and are not too effective in practice (Feldman, 2020; Paul et al., 2021). Active learning iteratively selects an informative subset of a larger pool of unlabeled data for annotation, which is ultimately used for supervised learning (Tharwat and Schenck, 2023; Ren et al., 2021; Beluch et al., 2018; Kirsch et al., 2019). Subsampling deletes instances of certain groups or classes when datasets are imbalanced, aiming to reduce bias and improve robustness of the downstream classifiers (Chawla et al., 2002; Barua et al., 2012; Chaudhuri et al., 2023).
@@ 2 Data Pruning & Fairness
More recently, data pruning emerged as a new research direction that simply removes parts of the dataset while maintaining strong model performance. In contrast to previous techniques, data pruning selects a subset of the original, fully labeled, and not necessarily imbalanced dataset, all while enjoying strong results in deep learning applications. Data pruning algorithms use the entire dataset $\mathcal{D}=\{X_{i},y_{i}\}_{i=1}^{N}$ to optimize a preliminary query model $\phi_{\theta}$ parameterized by $\theta$ that most often assigns “utility” scores $A(\phi,X)$ to each training sample $X$; then, the desired fraction $s$ of the least useful instances is pruned from the dataset, yielding a sparse subset $\mathcal{D}_{s}=\{X\colon A(\phi,X)\geq\text{quantile}[A(\phi,\mathcal{D}),s]\}$. In their seminal work, Toneva et al. (2019) let $A(\phi,X)$ be the number of times $(X,y)$ is both learned and forgotten while training the query model. Paul et al. (2021) design a “difficulty measure” $A(\phi,X)=\lVert\sigma[\phi(X)]-y\rVert_{2}$ where $\sigma$ denotes the softmax function and $y$ is one-hot. These scores, coined EL2N, are designed to approximate the GraNd metric $A(\phi_{\theta},X)=\lVert\nabla_{\theta}\mathcal{L}[\phi(X),y]\rVert_{2}$, which is simply the $\ell_{2}$ norm of the parameter gradient of the loss $\mathcal{L}$ computed at $(X,y)$. Both EL2N and GraNd scores require only a short training round (e.g., 10 epochs) of the query model. He et al. (2023) propose to select samples according to their dynamic uncertainty throughout training of the query model. For each training epoch $k$, they estimate the variance of the target probability $\sigma_{y}[\phi(X)]$ across a fixed window of $J$ previous epochs, and finally average those scores across $k$. Since CoreSet approaches are highly relevant for data pruning, we also consider one such label-agnostic procedure that greedily selects training samples that best jointly cover the data embeddings extracted from the penultimate layer of the trained query model $\psi$ (Sener and Savarese, 2018). While all these methods come from various contexts and with different motivations, several studies show that the scores computed by many of them exhibit high cross-correlation (Sorscher et al., 2022; Kwok et al., 2024).
Figure 1: The average test performance of various data pruning algorithms against dataset density (fraction of samples remaining after pruning) and measures of class-wise fairness. Top: VGG-16 on CIFAR-10, Center: VGG-19 on CIFAR-100, Bottom: ResNet-18 on TinyImageNet. All results averaged over $3$ random seeds. Error bands represent min/max.
Fairness & Evaluation.
@@
Fairness in machine learning concerns the issue of non-uniform accuracy over the data distribution (Caton and Haas, 2023). The majority of research focuses on group fairness where certain, often under-represented or sensitive, groups of the population enjoy worse predictive performance (Hashimoto et al., 2018; Thomas McCoy et al., 2020). In general, groups are subsets of classes, and worst-group accuracy is a standard optimization criterion in this domain (Sagawa* et al., 2020; Kirichenko et al., 2023; Rudner et al., 2024). The focus of this study is a special case of group fairness, classification bias, where groups are directly defined by class attributions. Classification bias commonly arises in the context of imbalanced datasets where tail classes require upsampling or importance weighting to produce models with strong worst-class performance (Cui et al., 2019; Barua et al., 2012; Tan et al., 2020; Chaudhuri et al., 2023). For balanced datasets, classification bias is found to be exacerbated by adversarial training (Li and Liu, 2023; Benz et al., 2021; Xu et al., 2021; Nanda et al., 2021; Ma et al., 2022) and network pruning (Paganini, 2020; Joseph et al., 2020; Tran et al., 2022; Good et al., 2022). These works measure classification bias in various ways ranging from visual inspection to confidence intervals for the slope of the linear regression that fits accuracy deviations around the mean across classes. Another common approach is to measure the class- or group-wise performance gaps relative to the baseline model (without adversarial training or pruning) (Hashemizadeh et al., 2023; Dai et al., 2023; Lin et al., 2022). However, this metric ignores the fact that this original model can be and often is highly unfair itself, and counts any potential improvement over it as unfavorable. Instead, we utilize a simpler and more natural suite of metrics to measure classification bias. Given accuracy (recall) $r_{k}$ for each class $k\in[K]$, we report (1) worst-class accuracy $\min_{k}r_{k}$, (2) difference between the maximum and the minimum recall $\max_{k}r_{k}$-$\min_{k}r_{k}$ (Joseph et al., 2020), and (3) standard deviation of recalls $\text{std}_{k}r_{k}$ (Ma et al., 2022).
Data Pruning Meets Fairness.
@@
Data Pruning is Currently Unfair.
@@
We conduct the first comprehensive comparative study of data pruning through the lens of class-wise fairness (Figure 1). First, we note that no pruning method is uniformly state-of-the-art, and the results vary considerably across model-dataset pairs. Among all algorithms, Dynamic Uncertainty with VGG-19 on CIFAR-100 presents a particularly interesting and instructive case. While it is arguably the best with respect to the average test performance, it fails miserably across all fairness metrics. Figure 2 reveals that it actually removes entire classes already at $10\%$ of CIFAR-100. Therefore, it seems plausible that Dynamic Uncertainty simply sacrifices the most difficult classes to retain strong performance on the easier ones. Figure 3 confirms our hypothesis: in contrast to other algorithms, at low density levels, Dynamic Uncertainty tends to prune classes with lower baseline accuracy (obtained from training on the full dataset) more aggressively, which entails a catastrophic classification bias hidden underneath a deceptively strong average accuracy. This observation presents a compelling argument for using criteria beyond average test performance for data pruning algorithms, particularly emphasizing classification bias.
@@
Overall, existing algorithms exhibit poor robustness to bias, although several of them improve ever slightly over the full dataset. In particular, EL2N and GraNd achieve a relatively low classification bias, closely followed by Forgetting. At the same time, Forgetting has a substantially stronger average test accuracy compared to these three methods, falling short of Random only after pruning more than $60\%$ from CIFAR-100. Forgetting produces more balanced datasets than EL2N and GraNd at low densities (Figure 2), and tends to prune “easier” classes more aggressively compared to all other methods (Figure 3). These two properties seem to be beneficial, especially when the available training data is scarce.
While the results in Figure 1 uncover strong classification bias, existing literature hints that data pruning should in fact be effective in mitigating the classification bias because it bears similarity to some fairness-driven optimization approaches. The majority of well-established and successful algorithms that optimize for worst-group performance apply importance weighting to effectively re-balance the long-tailed data distributions (cf. cost-sensitive learning (Elkan, 2001)). Taking this paradigm to an extreme, data pruning assigns zero weights to the pruned samples in the original dataset, trading some learning signal for faster convergence, and resource and memory efficiency. Furthermore, similarly to EL2N or Forgetting, the majority of cost-sensitive methods estimate class- or group-specific weights according to their difficulty. Thus, Sagawa* et al. (2020) optimize an approximate minimax DRO (distributionally robust) objective using a weighted sum of group-wise losses, putting higher mass on high-loss groups; Sinha et al. (2022) weigh samples by the current class-wise misclassification rates measured on a holdout validation set; Liu et al. (2021) pre-train a reference model to estimate the importance coefficients of groups for subsequent re-training. Similar cost-weighting strategies are adopted for robust knowledge distillation (Wang et al., 2023; Lukasik et al., 2022) and online batch selection (Kawaguchi and Lu, 2020; Mindermann et al., 2022). These techniques compute weights at the level of classes or sensitive groups and not for individual training samples. In the context of data pruning, this would correspond to selecting the appropriate density ratios across classes but subsampling randomly otherwise. While existing pruning algorithms automatically balance classes according to their validation errors fairly well (Figure 3), their overly scrupulous filtering within classes may be suboptimal. In support of this view, Ayed and Hayou (2023) formally show that integrating random sampling into score-based pruning procedures improves their average performance. In fact, random pruning is regarded as a notoriously strong baseline in active learning and data pruning (see Figure 1). In the next section, we propose and evaluate a random pruning algorithm that respects difficulty-based class-wise ratios.
3 Method & Results

### v2

# DRoP: Distributionally Robust Pruning

@@ Abstract
In the era of exceptionally data-hungry models, careful selection of the training data is essential to mitigate the extensive costs of deep learning. Data pruning offers a solution by removing redundant or uninformative samples from the dataset, which yields faster convergence and improved neural scaling laws. However, little is known about its impact on classification bias of the trained models. We conduct the first systematic study of this effect and reveal that existing data pruning algorithms can produce highly biased classifiers. We present theoretical analysis of the classification risk in a mixture of Gaussians to argue that choosing appropriate class pruning ratios, coupled with random pruning within classes has potential to improve worst-class performance. We thus propose DRoP, a distributionally robust approach to pruning and empirically demonstrate its performance on standard computer vision benchmarks. In sharp contrast to existing algorithms, our proposed method continues improving distributional robustness at a tolerable drop of average performance as we prune more from the datasets.
@@ 1 Introduction
The ever-increasing state-of-the-art performance of deep learning models requires exponentially larger volumes of training data according to neural scaling laws (Hestness et al., 2017; Kaplan et al., 2020; Rosenfeld et al., 2020; Gordon et al., 2021). However, not all collected data is equally important for learning as it contains noisy, repetitive, or uninformative samples. A recent research thread on data pruning is concerned with removing those unnecessary data, resulting in improved convergence speed, scaling, and resource efficiency (Toneva et al., 2019; Paul et al., 2021; He et al., 2023). These methods design scoring mechanisms to assess the utility of each sample, often measured by its difficulty or uncertainty as approximated during a preliminary training round, that guides pruning. Sorscher et al. (2022) report that selecting high-quality data using these techniques can trace a Pareto optimal frontier, beating the notorious power scaling laws; Kolossov et al. (2024) demonstrate that data selection can improve training.
@@
A recent study by (Pote et al., 2023) hints that data pruning may mitigate distributional bias in trained models, which is a well-established issue in AI systems concerning the performance disparity across classes (Dwork et al., 2012; Hardt et al., 2016b) or protected, minority groups of the population (e.g., race or gender). In Section 2, we conduct the first systematic evaluation of various pruning algorithms with respect to classification bias on a variety of benchmarks and conclude otherwise. For example, we find that Dynamic Uncertainty He et al. (2023) achieves superior average test performance with VGG-19 on CIFAR-100 and ResNet-50 on ImageNet but fails miserably in terms of worst-class accuracy. Thus, we argue that it is imperative to benchmark pruning algorithms using a more comprehensive suite of metrics that reflect classification bias, and to develop solutions that address distributional robustness directly.
@@
To understand the fundamental principles that exacerbate classification bias for various pruning methods, we present a theoretical analysis of linear decision rules for a mixture of two isotropic Gaussians in Section 3. It also illustrates how simple random subsampling with difficulty-aware class-wise pruning quotas may yield better worst-class performance compared to existing, finer pruning algorithms that operate on a sample level. Based on these observations, we design a first “robustness-aware” data pruning protocol, coined Distributionally Robust Pruning (DRoP). It selects target class proportions based on the corresponding class-wise error rates computed on a hold-out validation set after a preliminary training round on the full dataset. While these target quotas already improve robustness when combined with existing pruning algorithms, they particularly shine when applied together with random pruning in accordance with our theoretical analysis. Pruning with DRoP substantially reduces the classification bias of models even compared to training on the full dataset while offering enhanced data efficiency. We further verified the effectiveness of our method for initially imbalanced datasets and in the context of group robustness.
In hindsight, DRoP carries some analogies with distributionally robust optimization methods that address class/group bias, such as subsampling (Barua et al., 2012; Cui et al., 2019; Tan et al., 2020) or cost-sensitive learning that constructs error-based importance scores to reweigh the training loss, emphasizing data from the more difficult or minority classes (Sinha et al., 2022; Sagawa* et al., 2020; Liu et al., 2021; Idrissi et al., 2022; Wang et al., 2023; Lukasik et al., 2022; Chen et al., 2017). DRoP emulates this behavior with data pruning, retaining more samples from the more difficult classes. We discuss these similarities and compare DRoP to a prototypical cost-sensitive method, CDB-W (Sinha et al., 2022), on a variety of dataset types to further illustrate the effectiveness of our method.
@@
  • In Section 2, using a standard computer vision testbed, we conduct the first comprehensive evaluation of existing data pruning algorithms through the lens of classification bias for a variety of datasets and architectures;
  • In Section 3, we provide our theoretical analysis in the Gaussian mixture model to illustrate increased bias of current pruning methods, show how to optimize for worst-class risk and thus motivate our proposed solution;
  • In Section 4, we propose a random pruning procedure with error-based class ratios coined DRoP, and verify its effectiveness in drastically reducing the classification bias. We also provide ablations to illustrate the strength of DRoP as a pruning method, which in some range is even able to improve over full-dataset cost-sensitive learning.
2 Data Pruning & Distributional Robustness
In the era where the training corpora of contemporary models are of web-scale size, improving data efficiency has become the focus of practitioners and researchers alike. The corresponding literature is exceptionally rich, with a few fruitful and relevant research threads. Dataset distillation replaces the original samples with synthetically generated ones that bear compressed, albeit not as much interpretable, training signal (Sucholutsky & Schonlau, 2021; Cazenavette et al., 2022; Such et al., 2020; Zhao & Bilen, 2023; Nguyen et al., 2021; Feng et al., 2024). CoreSet methods select representative samples that jointly capture the entire data manifold (Sener & Savarese, 2018; Guo et al., 2022; Zheng et al., 2023; Agarwal et al., 2020; Mirzasoleiman et al., 2020; Welling, 2009); they yield weak generalization guarantees for non-convex problems and are not too effective in practice, especially on larger datasets (Feldman, 2020; Paul et al., 2021). Active learning iteratively selects an informative subset of a larger pool of unlabeled data for annotation, which is ultimately used for supervised learning (Tharwat & Schenck, 2023; Ren et al., 2021; Beluch et al., 2018; Kirsch et al., 2019). Subsampling deletes instances of certain groups or classes when datasets are imbalanced, aiming to reduce bias and improve robustness of the downstream classifiers (Chawla et al., 2002; Barua et al., 2012; Chaudhuri et al., 2023).
@@ 2 Data Pruning & Distributional Robustness
More recently, data pruning emerged as a new research direction that simply removes parts of the dataset while maintaining strong model performance. In contrast to previous techniques, data pruning selects a subset of the original, fully labeled, and not necessarily imbalanced dataset, all while enjoying strong results in deep learning applications. Data pruning algorithms use the entire dataset $\mathcal{D}=\{X_{i},y_{i}\}_{i=1}^{N}$ to optimize a preliminary query model $\psi_{\theta}$ parameterized by $\theta$ that most often assigns “utility” scores $A(\psi,X)$ to each training sample $X$; then, the desired fraction $s$ of the least useful instances is pruned from the dataset, yielding a sparse subset $\mathcal{D}_{s}=\{X\colon A(\psi,X)\geq\text{quantile}[A(\psi,\mathcal{D}),s]\}$. In their seminal work, Toneva et al. (2019) let $A(\psi,X)$ be the number of times $(X,y)$ is both learned and forgotten while training the query model. Paul et al. (2021) design a “difficulty” measure $A(\psi,X)=\lVert\sigma[\psi(X)]-y\rVert_{2}$ where $\sigma$ denotes the softmax function and $y$ is one-hot. These scores, coined EL2N, are designed to approximate the GraNd metric $A(\psi_{\theta},X)=\lVert\nabla_{\theta}\mathcal{L}[\psi(X),y]\rVert_{2}$, which is simply the $\ell_{2}$-norm of the parameter gradient of the loss $\mathcal{L}$ computed at $(X,y)$. Both EL2N and GraNd scores require only a short training round (e.g., 10 epochs) of the query model. He et al. (2023) propose to select samples according to their dynamic uncertainty throughout training of the query model. For each training epoch $k$, they estimate the variance of the target probability $\sigma_{y}[\psi(X)]$ across a fixed window of $J$ previous epochs, and finally average those scores across $k$. Since CoreSet approaches are highly relevant for data pruning, we also consider one such label-agnostic procedure that greedily selects training samples that best jointly cover the data embeddings extracted from the penultimate layer of the trained query model $\psi$ (Sener & Savarese, 2018). While all these methods come from various contexts and with different motivations, several studies show that the scores computed by many of them exhibit high cross-correlation (Sorscher et al., 2022; Kwok et al., 2024).
Robustness & Evaluation Metrics.
@@
Distributional robustness in machine learning concerns the issue of non-uniform accuracy over the data distribution (Sagawa* et al., 2020). The majority of the vast research in this area focuses on group robustness where certain, often under-represented or sensitive, groups of the population have worse predictive performance (Hashimoto et al., 2018; Thomas McCoy et al., 2020). In general, groups can be subsets of classes, and worst-group accuracy is a standard optimization criterion in this domain (Sagawa* et al., 2020; Kirichenko et al., 2023; Rudner et al., 2024). As a special case, fairness in machine learning aims to mitigate disparity of model performance across society and demographic groups and uses specific criteria such as equal opportunity and equalized odds (Hardt et al., 2016a; Caton & Haas, 2023). Many well-established algorithms in the space of distributional robustness apply importance weighting to effectively re-balance the long-tailed data distributions (cf. cost-sensitive learning (Elkan, 2001)). Thus, Sagawa* et al. (2020) optimize an approximate minimax DRO (distributionally robust optimization) objective using a weighted sum of group-wise losses, putting higher mass on high-loss groups; Sinha et al. (2022) weigh samples by the current class-wise misclassification rates measured on a holdout validation set; Liu et al. (2021) pre-train a reference model to estimate the importance factors of groups for subsequent re-training. Similar cost-weighting strategies are adopted for robust knowledge distillation (Wang et al., 2023; Lukasik et al., 2022) and online batch selection (Kawaguchi & Lu, 2020; Mindermann et al., 2022). Note that these techniques compute weights at the level of classes or sensitive groups and not for individual training samples. The focus of our study is a special case of group robustness, classification bias, where groups are directly defined by class attributions. Classification bias commonly arises in the context of imbalanced datasets where tail classes require upsampling or reweighting to produce models with strong worst-class performance (Cui et al., 2019; Barua et al., 2012; Tan et al., 2020; Chaudhuri et al., 2023). However, classification bias is also studied for balanced datasets, and is found to be exacerbated by adversarial training (Li & Liu, 2023; Benz et al., 2021; Xu et al., 2021; Nanda et al., 2021; Ma et al., 2022) and network parameter pruning (Paganini, 2020; Joseph et al., 2020; Tran et al., 2022; Good et al., 2022). Following mainstream prior work Li & Liu (2023); Zayed et al. (2023), we will use a natural suite of metrics to measure classification bias. Given accuracy (recall) $r_{k}$ for each class $k\in[K]$, we report (1) worst-class accuracy $\min_{k}r_{k}$, (2) difference between the maximum and the minimum recall, $\max_{k}r_{k}-\min_{k}r_{k}$ (Joseph et al., 2020), and (3) standard deviation of recalls $\text{std}_{k}r_{k}$ (Ma et al., 2022).
Data Pruning Meets Robustness.
@@
Figure 1: The average test performance of various data pruning algorithms against dataset density (fraction of samples remaining after pruning) and worst-class accuracy. All results averaged over $3$ random seeds. Error bands represent min/max. Additional plots can be found in Appendix C.
Figure 2: Pruning Exacerbates Bias: Dynamic Uncertainty applied to CIFAR-100. See Appendix E for similar plots for other pruning methods and models. Left: Sorted class densities at different dataset density levels. We also report the minimum number of samples per class (SPC) at $10\%$ dataset density. Right: Full dataset test class-wise accuracy against dataset density. We also report the correlation coefficient between these two quantities across classes, averaged over $5$ dataset densities.
Data Pruning Exacerbates Bias.
@@
Figure 1 illustrates our evaluation of data pruning through the lens of class-wise robustness for two vision benchmarks, CIFAR-100 and TinyImageNet (additional plots for other metrics can be found in Appendix C). We present the experimental details and hyperparameters in Appendix A. First, we note that no pruning method is uniformly state-of-the-art, and the results vary considerably across model-dataset pairs. Among all algorithms, Dynamic Uncertainty with VGG-19 on CIFAR-100 and ResNet-50 on ImageNet presents a particularly interesting and instructive case. While it is arguably the best with respect to the average test performance, it fails miserably across all robustness metrics. Figure 2 left reveals that it actually removes entire classes already at $10\%$ of CIFAR-100. Therefore, it seems plausible that Dynamic Uncertainty simply sacrifices the most difficult classes to retain strong performance on the easier ones. Figure 13 confirms our hypothesis: in contrast to other algorithms, at low density levels, Dynamic Uncertainty tends to prune classes with lower baseline accuracy (obtained from training on the full dataset) more aggressively, which entails a catastrophic classification bias hidden underneath a deceptively strong average accuracy. This observation presents a compelling argument for using criteria beyond average test performance for data pruning algorithms, particularly emphasizing classification bias.
@@
Overall, all studied algorithms exhibit poor robustness to bias, although several of them improve ever so slightly over the full dataset. In particular, EL2N and GraNd achieve a relatively low classification bias, closely followed by Forgetting. At the same time, Forgetting has a substantially stronger average test accuracy compared to these three methods, falling short of Random only after pruning more than $60\%$ from CIFAR-100. As seen from additional plots in Appendix E, Forgetting produces more balanced datasets than EL2N and GraNd at low densities (Figure 14), and tends to prune “easier” classes more aggressively compared to all other methods (Figure 13). These two properties seem to be beneficial, especially when the available training data is scarce.
3 Theoretical Analysis
In this section, we derive analytical results for data pruning in a toy model of binary classification for a mixture of two univariate Gaussians with linear classifiers. Perhaps surprisingly, in this framework we can derive worst-class optimal densities as well as demonstrate how prototypical pruning algorithms fail with respect to class robustness.
$R(t)=\phi_{0}R_{0}(t)+\phi_{1}R_{1}(t)$ (2)
Figure 3: The effect of different pruning procedures on the solution mixture of Gaussians problem with $\mu_{0}=-1$, $\mu_{1}=1$, $\sigma_{0}=0.5$, $\sigma_{1}=1$, and $\phi_{0}=\phi_{1}$. Pruning to dataset density $d=50\%$. Left: Random pruning with the optimal class-wise densities that satisfy $d_{1}\phi_{1}\sigma_{0}=d_{0}\phi_{0}\sigma_{1}$. Middle: SSP. Right: Random pruning with respect to class ratios provided by the SSP algorithm. All results averaged across $10$ datasets $\{D_{i}\}_{i=1}^{10}$ each with $400$ points. The average ERM is $\overline{T}=\frac{1}{10}\sum_{i=1}^{10}T(D^{\prime}_{i})$ fitted to pruned datasets $D^{\prime}_{i}$. The class risks of the average and worst-class optimal decisions for this Gaussian mixture are $R_{0}[t^{*}(1)]=4.8\%$, $R_{1}[t^{*}(1)]=12.1\%$, and $R_{0}(\hat{t})=R_{1}(\hat{t})=9.1\%$.
$\hat{t}=(\mu_{0}\sigma_{1}+\mu_{1}\sigma_{0})/(\sigma_{0}+\sigma_{1}).$ (4)
Note that $\hat{t}<t^{*}(1)$. This means that in the balanced case $\hat{t}$ is closer to $\mu_{0}$, the mean of the “easier” class. To understand how we should prune to achieve optimal worst-class accuracy, we compute class priors $\phi_{0}$ and $\phi_{1}$ that guarantee that the average risk minimization (Equation 2) achieves the best worst-class risk. These priors can be seen as the optimal class proportions in a “robustness-aware” dataset. Observe that, from Equation 3, $t^{*}(\sigma_{0}/\sigma_{1})=\hat{t}$ because the logarithm in the discriminant vanishes and we obtain Equation 4. Therefore, the optimal average risk minimizer coincides with the solution to the worst-case error over classes, and is achieved when the class priors $\phi_{0},\phi_{1}$ satisfy
$\phi_{0}/\phi_{1}=\sigma_{0}/\sigma_{1}.$ (5)
Intuitively, sufficiently large datasets sampled from $\mathcal{M}$ with class-conditionally independent samples mixed in proportions $N_{0}$ and $N_{1}$ should have their empirical 0-1 risk minimizers close to $t^{*}(N_{0}/N_{1})$. That is, randomly retaining $d_{k}$ fraction of samples in class $k=0,1$ where $d_{0}N_{0}\sigma_{1}=d_{1}N_{1}\sigma_{0}$ (see Equation 5) yields robust solutions with equal statistical risk across classes and, hence, optimal worst-class error. Figure 3 (left) confirms our analysis: the average empirical risk minimizer (ERM) $\overline{T}$ fitted to datasets pruned in this fashion lands near $\hat{t}$.
The class-conditional independence assumption we made above is crucial. While it is respected when subsampling randomly within each class, it clearly does not hold for existing, more sophisticated, data pruning algorithms. Therefore, even though they tend to prune easier classes more aggressively as evident from Figure 13, they rarely enjoy any improvement of the worst-class performance compared to the original dataset. We further illustrate this observation by replicating a supervised variant of the Self-Supervised Pruning (SSP) developed by Sorscher et al. (2022) for ImageNet. In this algorithm, we remove samples globally (i.e., irrespective of their class membership) located within a certain margin $M>0$ of their class means. Intuitively, this algorithm discards the easiest or the most representative samples from the most densely populated regions of the distribution. As illustrated in Figure 3 (middle), this method fortuitously prunes the easier class more aggressively; indeed, the amount of samples removed from a class with mean $\mu$ and variance $\sigma^{2}$ is proportional to the area under the probability density function over the pruning interval $[\mu-M,\mu+M]$, which is larger for smaller values of $\sigma$. Nevertheless, if the original dataset has classes mixed in proportions $N_{0}$ and $N_{1}$, the solution remains close to $t^{*}(N_{0}/N_{1})$ even after pruning, as we show formally in Appendix B.3. On the other hand, random subsampling according to class-wise pruning proportions defined by SSP does improve worst-class accuracy, as illustrated in the right plots of Figure 3. This corresponds to our observation in Figure 5 that random pruning respecting the class proportions discovered by GraNd and Forgetting often improves robustness compared to these methods themselves.
Figure 4: (a): Class-wise risk ratios of the optimal solution $t^{*}=t^{*}(\phi_{0}/\phi_{1})$ vs. optimal ratios based on Equation 5 computed for various $\sigma_{0}<\sigma_{1}$ drawn uniformly from $[10^{-2},10^{2}]$ and $\phi_{0}\sim U[0,1]$ and $\phi_{1}=1-\phi_{0}$. The results are independent of $\mu_{0},\mu_{1}$. (b): Random pruning with DRoP. Left: $d=75\%$; Right: $d=50\%$.
A priori, it is unclear how exactly the variance-based worst-case optimal pruning quotas in Equation 5 generalize to deep learning, and in particular how they can lead to practical pruning algorithms. One could connect our theory to the feature distribution in the penultimate layer of neural networks and determine class densities from cluster variances around class means. In fact, SSP (Sorscher et al., 2022) uses such metrics to determine pruning scores for samples, using a pretrained model. However, such variances are noisy, especially for smaller class sizes, and hard to connect to class accuracies. Therefore, instead, we propose to use class errors as a proxy for these variances in our DRoP method introduced in Section 4. More formally, given a dataset originally with class sizes $N_{0}$ and $N_{1}$, we replace the worst-class optimal condition $d_{0}N_{0}\sigma_{1}=d_{1}N_{1}\sigma_{0}$ with $d_{0}R_{1}[t^{*}(N_{0}/N_{1})]=d_{1}R_{0}[t^{*}(N_{0}/N_{1})]$. To motivate this practical approach, Figure 4a shows that these new error-based class densities approximate the optimal variance-based ones fairly well, especially when $\sigma_{1}/\sigma_{0}$ is small. Figure 4b demonstrates that random pruning according to DRoP class proportions lands the average ERM near the worst-class optimal value. Thus, even though error-based class quotas do not enjoy simple theoretical guarantees, they still operate near-optimally in this toy setup.
4 DRoP: Distributionally Robust Pruning